Write down the equation for the alpha decay of Gold-185. (The symbol for gold is Au and its atomic number is 79). What is the mass of the daughter nuclide in your equation?
1
Expert's answer
2020-03-23T10:15:03-0400

Solution:

The equation for the alpha decay of Gold-185:




alpha particle = (4/2)He = Helium-4.

Ir-181 = Iridium-181 - daughter nuclide in this equation.

181 - is the mass of the daughter nuclide (Ir).


Answer:

The equation for the alpha decay: (185/79)Au = (181/77)Ir + (4/2)He;

181 is the mass of the daughter nuclide - (181/77)Ir.